OpenStudy (anonymous):

You (or your parents) plan to pay $1,275.00/month for mortgage. How much is the minimum realized income per month to the nearest penny?

OpenStudy (anonymous):

Would I form it like 28/100 = 1,275.00/x?

OpenStudy (adamk):

3215.17 - 1275.00 = 1940.19 for food, utilities, internet connection, clothes, etc

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

I have no idea how you got 3215.17, or anything.